OpenStudy (anonymous):

Solve. -17 - (3y - 1) = 2(y + 1) + 4y

OpenStudy (anonymous):

-17-3y+1=2y+2+4y -16-3y=6y+2 -16=9y+2 -18=9y Y=-2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem:)